AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 00:42:55 2025
Author
*
:
Subject
*
:
<p>10:30</p> <p>Merger OK, Tape server OK</p> <p>ADC data item statistics looks as before. OK</p> <p>FEE temperatures look OK</p> <p>System checks OK</p> <p>Leakage current stable</p> <p> </p> <p>10.48 FEE64 temperatures OK - attachment 1</p> <p>WR timestamps OK - attachment 2</p> <p>ADC data item stats - attachment 3</p> <p>per FEE64 Rate spectra - attachment 4</p> <p> why do we observe HEC data in aida15 asic #1?</p> <p>per FEE64 1.8.L spectra - attachments 5-6</p> <p> ADC to 20MeV FSR - not clear when these spwectra were last zero;d</p> <p> currently 750keV/u D + no target ?</p> <p>10.53 All histograms & stats zero'd</p> <p>Merger, Taper Server etc - attachments 7-8</p> <p> bind error 98 for NewMerger server - not clear when these messahes were generated - added some returns to monitor whether these, or other, messages are repeated</p> <p>DSSSD bias & leakage current OK - attachment 9</p> <p>11:30</p> <p>We noticed some odd peak structure in the rate energy histograms for the right hand side detectors. We took away the beam but this structure persisted. </p> <p>This was not observed in the total energy histograms, the reason for this is the applied multiplicity cut to the total histograms. Removing this cut the peaks are clearly visible (attachment 10). Putting the cut back in we see the peaks removed (attachment 11). These events correspond to some noisy strips on the detectors on the right hand side (see attachment 12). Notice the much larger rate on the RHS. Current ADC data items statistics show FEE modules for the RHS also much larger (attachment 13). We did not observe this with no beam (https://elog.ph.ed.ac.uk/CARME/554) but did with beam last night (https://elog.ph.ed.ac.uk/CARME/558) so there nust be some effect of having beam in the ring increasing the number of events on the RHS. Note the detectors are fully out and the events are localised in these strips. </p> <p>Note we have moved the detectorsand removed beam from the ring in this time period to investigate these peaks.</p> <p>12:50</p> <p>Run 10 continues. Temps OK. System checks OK. Merger OK, Tape server OK.</p> <p>Leakage current OK.</p> <p>Notice the 'beam peak' at 1.5 MeV is larger in the left hand side (attachment 14)</p> <p>14:26</p> <p>We have started the new_optimeren.py on pi@gas-stepper to try and align the gas target. Valves to the interaction chamber are closed.</p> <p>Setting the nozzle X=-13000, Y=1000 we observe an increase in E4 from 4E-10 mbar to 2E-9 mbar. Moving either side we observe a drop in E4. We take this as the gas target being aligned.</p> <p>Anton will now open the valves to the gas target.</p> <p>Run 10 stopped. </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown